Draw BMD and SFD for an overhang beam ABC with span BC overhanging. The
loading and span data is as follows
Span AB = 6 m; Full Span UDL = 20 kN/m,
Span BC = 2 m; Full Span UDL = 30 kN/m
Span BC = 6 m; Point Load at C = 15 kN.
1
Expert's answer
2021-08-31T01:49:02-0400



=-(15x8)+(-30x2)(7))+RBx6-(20x6)x3)

 =-120-420+6RB-360

 0=6RB-900 

900=6R 

RB=150  

RA=45  

Bending moments 

BMc=0 

BMb=-30-60=-90

BMA= -15x8-(6x70)+(150x6)-(120x3) 

=0 





THANK.,